Mr chen deposit a certain sum of money in a bank in the interest rate of the bank decreases from 3.75% per annum to 3.5% per annum. Mr chen's interest will decrease by $50 in a year. find the sum of money he deposites
Found 2 solutions by mananth, greenestamps:
Answer by mananth(16946)

Mr chen deposit a certain sum of money in a bank in the interest rate of the bank decreases from 3.75% per annum to 3.5% per annum. Mr chen's interest will decrease by $50 in a year. find the sum of money he deposites
Mr chen deposit a certain sum of money in a bank let it be x
interest

At 3.75% rate he gets 3.75%x
At 3.50% rate he gets 3.50%x
3.75%x-3.50%x =50
0.0375x-0.0350x =50
0.025x=50
x=50/0.025
x=2000
sum of money in a bank deposited = $2000



Answer by greenestamps(13200)



The other tutor who responded has a habit of supplying answers without checking to see if they are correct. A simple arithmetic error in her response leads to an incorrect answer.

The interest rate decreased by 1/4 of 1 percent, resulting in a decrease of $50 in interest.

A decrease of 1/4 of 1 percent resulting in a decrease of $50 in interest means that a decrease of 1 percent would result in a decrease of 4($50) = $200; that means the amount invested was 100($200) = $20,000.

ANSWER: $20,000